Solve
(D^2+DD'-6D'^2)z= ycosx
1
Expert's answer
2020-10-25T19:14:45-0400

Given,


"(D^2+DD'-6D'^2)z= y\cos x"

Now, Auxiliary equation will be


"m^2+m-6=0\\
\implies (m-2)(m+3)=0\\
\implies m=2,-3"

Hence,

"C.F=\phi_1(y+2x)+\phi_2(y-3x)"

Now,


"P.I=\frac{1}{(D^2+DD'-6D'^2)}y\cos x\\
\implies P.I=\frac{1}{(D+3D')(D-2D')}y\cos x\\
=\frac{1}{D+3D'}\int_{D-2D'}y\cos x \, dx\\
=\frac{1}{D+3D'}[(c_1-2x)\sin x-2\cos x]\\
=\frac{1}{D+3D'}(y\sin x-2\cos x)\\
=\int_{D+3D'}y\sin xdx-\int_{D+3D'}2\cos x dx\\
=\int_{D+3D'}(c_2+3x)\sin xdx-\int_{D+3D'}2\cos x dx\hspace{1cm}(y=c_2+3x)\\"


"\implies P.I=-y\cos x+\int\cos x \, dx =-y\cos x+\sin x"

Therefore the complete solution is

"C.F+P.I=\phi_1(y+2x)+\phi_2(y-3x)+-y\cos x+\sin x"